OpenStudy (anonymous):

15) Reduce the following rational expression to lowest terms, if possible. x^2 - 9 x + 14divide by x^2 - x - 2

OpenStudy (anonymous):

x^2-9x+14=(x-2)(x-7) x^2-x-2=(x-2)(x+1) divide the two and x-2 cancels (x-7)/(x+1)